OpenStudy (anonymous):

from a pack of 52 cards,a card is drawn at random.find the probablity that it is either an ace of red colour or a card of black colour.

OpenStudy (mathstudent55):

How many red aces are there in a deck of cards? How many black cards are there in a deck of cards?

OpenStudy (anonymous):

2 red aces and 26 black cards

OpenStudy (mathstudent55):

Correct. Since these two events are mutually exclusive, there are 2 + 26 = 28 desired outcomes out of a total of 52 possible outcomes. What is the probability of getting a desired outcome?

OpenStudy (anonymous):

28/52

OpenStudy (mathstudent55):

Correct. Just reduce the fraction.

OpenStudy (anonymous):

thank you very much

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

but why did you add them

OpenStudy (mathstudent55):

Since the desired outcome is 2 red aces OR a black card, you add the outcomes.

OpenStudy (anonymous):

because P(A or B) = P(A) + P(B) - P(A and B), but P(A and B) = 0,

OpenStudy (anonymous):

thank you to both mathstudent55 and sourwing

OpenStudy (mathstudent55):

P(A and B) = 0 because these two events are mutually exclusive. If you pick a red ace, it cannot possibly be a black card.

OpenStudy (mathstudent55):

wlcm